OpenStudy (anonymous):

solve and show work please :/ 1/3+1/2=1/x

OpenStudy (anonymous):

add on the left, then take the reciprocal to get \(x\)

OpenStudy (anonymous):

what ? ...

OpenStudy (anonymous):

what is \[\frac{1}{3}+\frac{1}{2}\]?

OpenStudy (anonymous):

5/6

OpenStudy (anonymous):

oh .. i was thinking about it too much -___- ive done way too much math today .. thanks !

hero (hero):

\(\large\frac{1}{3} + \frac{1}{2} = \frac{1}{x}\) Multiply the first fraction by 2/2; Multiply the second fraction by 3/3 to get: \(\large\frac{2}{6} + \frac{3}{6} = \frac{1}{x}\) Combine fractions on the left side to get: \(\large\frac{5}{6} = \frac{1}{x}\) Cross Multiply: \(5x = 6\) Divide both sides by 5 to get: \(\large x = \frac{6}{5}\)